Plasma; Floating Head Switch Not Working

If you have a question about the software please ask it here.

Plasma; Floating Head Switch Not Working

Postby Rad » Tue Jan 09, 2018 1:25 pm

Hi, I 'm using UCCNC + UC100 for CNC Plasma. The problem is that my Zaxis floating head limit switch is not working, I cannot see it being active on the diagnostic window of UCCNC when the switch is triggered via the floating head.

As a result of this, when doing initial height sensing before each pierce the torch would go down to the material close the limit switch via the float head but becuase the system does not know when the switch is triggered it continues diving into the material trying to find the surface of the material and crashes the torch into the material. This problem would happen occasionally but after a few tries of closing and re-opening UCCNC the Initial Height Sensing would function as it should.... But now it is Not working period. I tried replacing the limit switch with a new one and even tried shorting the terminals directly on the breakout board but I can still not see the switch being active on the diagnostic window of UCCNC. I am using pin 12 input on the breakout board for Z home.

My question is: is this problem UC100 related or is it a problem with my breakout board? FYI: all my other inputs for my THC are working fine on the breakout board.

Any help would be appreciated
Rad
 
Posts: 14
Joined: Fri Sep 15, 2017 12:10 pm

Re: Plasma; Floating Head Switch Not Working

Postby cncdrive » Tue Jan 09, 2018 1:54 pm

It is probably electrical issue.
If you do not see the input LED triggering then the UC100 does not receive the signal.
And if this is the case then the issue will be probably with the breakout board. If it has optocouplers then maybe the input LED is not driven with enough current to reliably drive out the opto.

To give you an example a HCPL2631 optocoupler requires 6.3mA of current to be reliably switch.
It has a 1.8V input forward voltage rating.
So, for example if the input is driven with 5Volts then the maximum series resistance can be R=(5V-1.8V)/0.0063A=508 Ohms, that will drive 6.3mAmps of current into the LED of the opto at 5Volts voltage.
If the series resistor is larger then it will not drive out the optocoupler properly.
Ofcourse this was just an example, as I do not know the circuitry of your BOB, so I do not know what optocoupler and what circuit is it using on the inputs.

And you can try to remove the breakout board from the UC100 and connect the input 12. to pins 18.-25. (all are grounds).
If it is still not changing when you short the mentioned pins then the input on the UC100 damaged.
Otherwise, if it works then the issue is with the BOB as described above.
cncdrive
Site Admin
 
Posts: 2190
Joined: Tue Aug 12, 2014 11:17 pm

Re: Plasma; Floating Head Switch Not Working

Postby Rad » Tue Jan 09, 2018 2:23 pm

Thanks, will connect the pins directly on the UC100 and see if it lights up on the diagnostic screen
Rad
 
Posts: 14
Joined: Fri Sep 15, 2017 12:10 pm


Return to Ask a question from support here

Who is online

Users browsing this forum: No registered users and 4 guests